OpenStudy (anonymous):

Solve log3 27 – x = 5

OpenStudy (mertsj):

\[\log_{3}(27-x)=5 \]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

\[3^5=27-x\]

OpenStudy (mertsj):

Solve that

OpenStudy (anonymous):

\[\log_{3} (27-x)= 5\] \[\log(27-x)/\log3 = 5\] \[\log(27-x)= 5\log3\] \[\log(27-x) = \log3^{5}\] \[27 - x = 3^{5} ##canceling log by taking exponential both sides##\]